eBook Calculator Single Plantwide Factory Overhead Rate Salty Sensations Snacks Company manufactures three types of snack fooTotal Factory Overhead Per Case Factory Overhead Tortilla chips $ Potato chips Pretzels Total
Answer #1

1) Budgeted hours = (4000*.20+5000*.15+2500*.40) = 2550 Hour

Predetermine overhead rate = 153000/2550 = 60 per hour

2) Calculate following

Total factory overhead Per Case factory overhead
Tortilla chips 800*60 = 48000 60*.20 = 12
Potato chips 750*60 = 45000 60*.15 = 9
Pretzels 60000 60*.40= 24
Total 153000